# ParallelTime: Dynamically Weighting the Balance of Short- and Long-Term Temporal Dependencies

### v1

@@ Abstract
Modern multivariate time series forecasting primarily relies on two architectures: the Transformer with attention mechanism and Mamba. In natural language processing, an approach has been used that combines local window attention for capturing short-term dependencies and Mamba for capturing long-term dependencies, with their outputs averaged to assign equal weight to both. We find that for time-series forecasting tasks, assigning equal weight to long-term and short-term dependencies is not optimal. To mitigate this, we propose a dynamic weighting mechanism, ParallelTime Weighter, which calculates interdependent weights for long-term and short-term dependencies for each token based on the input and the model’s knowledge. Furthermore, we introduce the ParallelTime architecture, which incorporates the ParallelTime Weighter mechanism to deliver state-of-the-art performance across diverse benchmarks. Our architecture demonstrates robustness, achieves lower FLOPs, requires fewer parameters, scales effectively to longer prediction horizons, and significantly outperforms existing methods. These advances highlight a promising path for future developments of parallel Attention-Mamba in time series forecasting. The implementation is readily available at: ParallelTime GitHub .
@@ 1 Introduction
  • We demonstrate that parallel Mamba-Attention architecture is currently the most effective approach for long-term time series forecasting, highlighting a promising direction for future advancements in time series forecasting models.
@@ 2 Related Work
Transformer Vaswani et al. (2017), leveraging causal self-attention layers and feed-forward networks, has laid a powerful foundation for time series forecasting (Zhou et al., 2021, 2022). A standout example is PatchTST (Nie et al., 2023), which achieves SOTA performance by utilizing channel independence to process each variable in a multivariate time series separately. By feeding contiguous time series patches as tokens into a standard self-attention mechanism, PatchTST outperforms many previous models. In standard self-attention, each token attends to all preceding tokens to capture global dependencies. To focus on local patterns, windowed attention variants, such as those in LongFormer (Beltagy et al., 2020a) and Swin Transformer (Liu et al., 2021), restrict each token to attend only to the most recent $\mathbf{S}$ tokens, as illustrated in Figure 1 (B).
@@ 3 ParallelTime
Figure 3: The architecture of ParallelTime. The input time series 1 is sliced into non-overlapping patches. 2 Each patch is embedded and augmented with positional encoding. The resulting tokens are processed through $N$ stacked ParallelTime blocks. Each block first normalizes the input, applies Mamba computation and 3 windowed attention with a register in parallel, 4 weights their outputs using a ParallelTime Weighter mechanism, and then applies normalization followed by a nonlinear feedforward layer. Finally, the output is normalized 5 and processed through an expand-compress-projection mechanism to generate the horizon prediction.
@@ 3.3 Mamba and Windowed Attention with Global Registers
To incorporate global context, we introduce global register tokens, denoted as $\mathbf{W}_{\text{reg}}\in\mathbb{R}^{R\times dim}$, where $R$ is the number of registers and $dim$ is the embedding dimension. These tokens serve as a compact repository of domain-specific global information, providing the model with access to broader contextual cues. The input sequence $\mathbf{x}_{d}\in\mathbb{R}^{P\times dim}$, is concatenated with the global registers to form: $\mathbf{x}_{\text{cat}}=\text{Concat}(\mathbf{W}_{\text{reg}},\mathbf{x}_{d})\in\mathbb{R}^{(R+P)\times dim}$. This concatenated sequence is then processed by the causal multi-head windowed attention mechanism, yielding:
@@ 3.4 ParallelTime Weighter
3.5 Loss and Optimizer
We employed the Huber loss function (Huber, 1992; Wen et al., 2019), chosen for its enhanced robustness to outliers and contribution to improved training stability. The Huber loss is defined as:
$\mathcal{L}_{\text{huber}}\left(x_{t},\hat{x}_{t}\right)=\begin{cases}\left(x_{t}-\hat{x}_{t}\right)^{2}\frac{1}{2},&\text{if }\left|x_{t}-\hat{x}_{t}\right|\leq\delta.\\ \delta\times\left(\left|x_{t}-\hat{x}_{t}\right|-\frac{1}{2}\times\delta\right),&\text{else}.\end{cases}$ (1)
where $x_{t}$ is the true value, $\hat{x}_{t}$ is the predicted value, and $\delta$ is a hyperparameter modulating the balance between L1 and L2 loss characteristics.
For optimization, we used the Adam optimizer (Kingma and Ba, 2015), which provides efficient adaptive learning rate adjustments. Further details regarding the training process and specific hyperparameter settings can be found in Appendix 8.2.
@@ 4.1 Baselines and Experimental Setup
To assess the performance of our proposed ParallelTime, we compare it against several SOTA models for long time series forecasting. These include Transformer-based models such as PatchTST (Nie et al., 2023) and FEDFormer (Zhou et al., 2022), Linear models like DLinear (Zeng et al., 2023) and N-BEATS (Oreshkin et al., 2020), foundational models including Moment (Goswami et al., 2024), GPT4TS (Zhou et al., 2023), and TimesNet (Wu et al., 2023), as well as ensemble models LightTS (Campos et al., 2023) and Stationary. We evaluate all models on eight widely used datasets: Electricity, Weather, Illness, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). For detailed dataset descriptions, see Appendix  8.1.
@@
Methods ParallelTime MOMENT GPT4TS PatchTST DLinear TimesNet FEDFormer Stationary LightTS N-BEATS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.189 0.154 0.209 0.162 0.212 0.149 0.198 0.176 0.237 0.172 0.220 0.217 0.296 0.173 0.223 0.182 0.242 0.152 0.210
192 0.189 0.232 0.197 0.248 0.204 0.248 0.194 0.241 0.220 0.282 0.219 0.261 0.276 0.336 0.245 0.285 0.227 0.287 0.199 0.260
336 0.242 0.273 0.246 0.285 0.254 0.286 0.245 0.282 0.265 0.319 0.280 0.306 0.339 0.380 0.321 0.338 0.282 0.334 0.258 0.311
720 0.323 0.331 0.315 0.336 0.326 0.337 0.314 0.334 0.333 0.362 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.386 0.331 0.359
ETTh1 96 0.365 0.398 0.387 0.410 0.376 0.397 0.370 0.399 0.375 0.399 0.384 0.402 0.376 0.419 0.513 0.491 0.424 0.432 0.399 0.428
192 0.399 0.415 0.410 0.426 0.416 0.418 0.413 0.421 0.405 0.416 0.436 0.429 0.420 0.448 0.534 0.504 0.475 0.462 0.451 0.464
336 0.385 0.414 0.422 0.437 0.442 0.433 0.422 0.436 0.439 0.443 0.491 0.469 0.459 0.465 0.588 0.535 0.518 0.488 0.498 0.500
720 0.420 0.443 0.454 0.472 0.477 0.456 0.447 0.466 0.472 0.490 0.521 0.500 0.506 0.507 0.643 0.616 0.547 0.533 0.608 0.573
ETTh2 96 0.262 0.328 0.288 0.345 0.285 0.342 0.274 0.336 0.289 0.353 0.340 0.374 0.358 0.397 0.476 0.458 0.397 0.437 0.327 0.387
192 0.322 0.368 0.349 0.386 0.354 0.389 0.339 0.379 0.383 0.418 0.402 0.414 0.429 0.439 0.512 0.493 0.520 0.504 0.400 0.435
336 0.312 0.370 0.369 0.408 0.373 0.407 0.329 0.380 0.448 0.465 0.452 0.452 0.496 0.487 0.552 0.551 0.626 0.559 0.747 0.599
720 0.399 0.434 0.403 0.439 0.406 0.441 0.379 0.422 0.605 0.551 0.462 0.468 0.463 0.474 0.562 0.560 0.863 0.672 1.454 0.847
ETTm1 96 0.284 0.337 0.293 0.349 0.292 0.346 0.290 0.342 0.299 0.343 0.338 0.375 0.379 0.419 0.386 0.398 0.374 0.400 0.318 0.367
192 0.329 0.366 0.326 0.368 0.332 0.372 0.332 0.369 0.335 0.365 0.374 0.387 0.426 0.441 0.459 0.444 0.400 0.407 0.355 0.391
336 0.365 0.391 0.352 0.384 0.366 0.394 0.366 0.392 0.369 0.386 0.410 0.411 0.445 0.459 0.495 0.464 0.438 0.438 0.401 0.419
720 0.424 0.430 0.405 0.416 0.417 0.421 0.416 0.420 0.425 0.421 0.478 0.450 0.543 0.490 0.585 0.516 0.527 0.502 0.448 0.448
ETTm2 96 0.162 0.252 0.170 0.260 0.173 0.262 0.165 0.255 0.167 0.269 0.187 0.267 0.203 0.287 0.192 0.274 0.209 0.308 0.197 0.271
192 0.218 0.291 0.227 0.297 0.229 0.301 0.220 0.292 0.224 0.303 0.249 0.309 0.269 0.328 0.280 0.339 0.311 0.382 0.285 0.328
336 0.276 0.327 0.275 0.328 0.286 0.341 0.274 0.329 0.281 0.342 0.321 0.351 0.325 0.366 0.334 0.361 0.442 0.466 0.338 0.366
720 0.356 0.380 0.363 0.387 0.378 0.401 0.362 0.385 0.397 0.421 0.408 0.403 0.421 0.415 0.417 0.413 0.675 0.587 0.395 0.419
Illness 24 1.166 0.657 2.728 1.114 2.063 0.881 1.319 0.754 2.215 1.081 2.317 0.934 3.228 1.260 2.294 0.945 8.313 2.144 4.539 1.528
36 1.293 0.727 2.669 1.092 1.868 0.892 1.430 0.834 1.963 0.963 1.972 0.920 2.679 1.080 1.825 0.848 6.631 1.902 4.628 1.534
48 1.399 0.772 2.728 1.098 1.790 0.884 1.553 0.815 2.130 1.024 2.238 0.940 2.622 1.078 2.010 0.900 7.299 1.982 4.957 1.585
60 1.615 0.844 2.883 1.126 1.979 0.957 1.470 0.788 2.368 1.096 2.027 0.928 2.857 1.157 2.178 0.963 7.283 1.985 5.429 1.661
Electricity 96 0.128 0.222 0.136 0.233 0.139 0.238 0.129 0.222 0.140 0.237 0.168 0.272 0.193 0.308 0.169 0.273 0.207 0.307 0.131 0.228
192 0.148 0.240 0.152 0.247 0.153 0.251 0.157 0.240 0.153 0.249 0.184 0.289 0.201 0.315 0.182 0.286 0.213 0.316 0.153 0.248
336 0.163 0.258 0.167 0.264 0.169 0.266 0.163 0.259 0.169 0.267 0.198 0.300 0.214 0.329 0.200 0.304 0.230 0.333 0.170 0.267
720 0.197 0.288 0.205 0.295 0.206 0.297 0.197 0.290 0.203 0.301 0.220 0.320 0.246 0.355 0.222 0.321 0.265 0.360 0.208 0.298
Traffic 96 0.349 0.231 0.391 0.282 0.388 0.282 0.360 0.249 0.410 0.282 0.593 0.321 0.587 0.366 0.612 0.338 0.615 0.391 0.375 0.259
192 0.371 0.240 0.404 0.287 0.407 0.290 0.379 0.256 0.423 0.287 0.617 0.336 0.604 0.373 0.613 0.340 0.601 0.382 0.403 0.274
336 0.388 0.250 0.414 0.292 0.412 0.294 0.392 0.264 0.436 0.296 0.629 0.336 0.621 0.383 0.618 0.328 0.613 0.386 0.426 0.285
720 0.429 0.274 0.450 0.310 0.450 0.312 0.432 0.286 0.466 0.315 0.640 0.350 0.626 0.382 0.653 0.355 0.658 0.407 0.508 0.335
@@
Table 1: Full results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. The results obtained from (Goswami et al., 2024). Red: the best, Underline: the 2nd best.
MSE MAE Fwd FLOPs Fwd+Bwd FLOPs #Params
Pred Len ParallelTime PatchTST ParallelTime PatchTST ParallelTime PatchTST ParallelTime PatchTST ParallelTime PatchTST
96 0.349 (↓3.1%) 0.360 0.231 (↓7.2%) 0.249 8.41G (↓35%) 13.1G 25.2G (↓36%) 39.5G 614k (↓48%) 1194k
192 0.371 (↓2.1%) 0.379 0.240 (↓6.3%) 0.256 8.42G (↓37%) 13.5G 25.2G (↓37%) 40.5G 651k (↓67%) 1980k
336 0.388 (↓1.0%) 0.392 0.250 (↓5.3%) 0.264 8.45G (↓39%) 14.0G 25.3G (↓39%) 42.1G 707k (↓77%) 3160k
720 0.429 (↓0.7%) 0.432 0.274 (↓4.2%) 0.286 8.51G (↓44%) 15.4G 25.5G (↓44%) 46.3G 855k (↓86%) 6306k
@@
Table 2: Comparison of ParallelTime and PatchTST on the Traffic dataset. The table reports MSE, MAE, forward (Fwd) FLOPs (i.e., inference FLOPs), forward and backward (Fwd+Bwd) FLOPs (i.e., training FLOPs), and the number of parameters (#Params). Bold values indicate superior performance. ↓ indicates that lower values are better. The improvement percentages for ParallelTime over PatchTST are shown in parentheses.
@@ 4.2 Main Results
Comprehensive forecasting results are listed in Table 1, with the best performance highlighted in red and the second best underlined. A lower Mean Squared Error (MSE) and Mean Absolute Error (MAE) indicate more accurate predictions. Our proposed model, ParallelTime, demonstrates exceptional performance across a diverse set of datasets and prediction horizons, consistently outperforming a range of state-of-the-art models, ParallelTime achieves the best forecasting accuracy in a significant number of scenarios, particularly excelling in datasets such as Weather, ETTh1, ETTh2, ETTm2, Electricity, Traffic, and Illness.

### v2

@@ Abstract
Modern multivariate time series forecasting primarily relies on two architectures: the Transformer with attention mechanism and Mamba. In natural language processing, an approach has been used that combines local window attention for capturing short-term dependencies and Mamba for capturing long-term dependencies, with their outputs averaged to assign equal weight to both. We find that for time-series forecasting tasks, assigning equal weight to long-term and short-term dependencies is not optimal. To mitigate this, we propose a dynamic weighting mechanism, ParallelTime Weighter, which calculates interdependent weights for long-term and short-term dependencies for each token based on the input and the model’s knowledge. Furthermore, we introduce the ParallelTime architecture, which incorporates the ParallelTime Weighter mechanism to deliver state-of-the-art performance across diverse benchmarks. Our architecture demonstrates robustness, achieves lower FLOPs, requires fewer parameters, scales effectively to longer prediction horizons, and significantly outperforms existing methods. These advances highlight a promising path for future developments of parallel Attention-Mamba in time series forecasting. The implementation is readily available at: GitHub.
@@ 1 Introduction
  • We demonstrate that the parallel Mamba-Attention architecture is the most effective approach for long-term time series forecasting.
@@ 2 Related Work
Transformer Vaswani et al. (2017), leveraging causal self-attention layers and feed-forward networks, has laid a powerful foundation for time series forecasting (Zhou et al., 2021, 2022). A standout example is PatchTST (Nie et al., 2023), which achieves SOTA performance by utilizing channel independence to process each variable in a multivariate time series separately. By feeding contiguous time series patches as tokens into a standard self-attention mechanism, PatchTST outperforms many previous models. In standard self-attention, each token attends to all preceding tokens to capture global dependencies. To focus on local patterns, windowed attention variants, such as those in LongFormer (Beltagy et al., 2020a) and Swin Transformer (Liu et al., 2021), restrict each token to attend only to the most recent S{S} tokens, as illustrated in Figure 1 (B).
@@ 3 ParallelTime
Figure 3: The architecture of ParallelTime. The input time series 1 is sliced into non-overlapping patches. 2 Each patch is embedded and augmented with positional encoding. The resulting tokens are processed through NN stacked ParallelTime blocks. Each block first normalizes the input, applies Mamba computation and 3 windowed attention with a register in parallel, 4 weights their outputs using a ParallelTime Weighter mechanism, and then applies normalization followed by a nonlinear feedforward layer. Finally, the output is normalized 5 and processed through an expand-compress-projection mechanism to generate the horizon prediction.
@@ 3.3 Mamba and Windowed Attention with Global Registers
To incorporate global context, we introduce global register tokens, denoted as $\mathbf{W}_{\text{reg}}\in\mathbb{R}^{R\times dim}$, where $R$ is the number of registers and dimdim is the embedding dimension. These tokens serve as a compact repository of domain-specific global information, providing the model with access to broader contextual cues. The input sequence $\mathbf{x}_{d}\in\mathbb{R}^{P\times dim}$, is concatenated with the global registers to form: $\mathbf{x}_{\text{cat}}=\text{Concat}(\mathbf{W}_{\text{reg}},\mathbf{x}_{d})\in\mathbb{R}^{(R+P)\times dim}$. This concatenated sequence is then processed by the causal multi-head windowed attention mechanism, yielding:
@@ 4.1 Baselines and Experimental Setup
To assess the performance of our proposed ParallelTime, we compare it against several SOTA models for long time series forecasting. These include Transformer-based models such as PatchTST (Nie et al., 2023) iTransformer (Liu et al., 2024) and FEDFormer (Zhou et al., 2022), Mamba models S-Mamba (Wang et al., 2024), Linear model DLinear (Zeng et al., 2023), foundational models including Moment (Goswami et al., 2024), GPT4TS (Zhou et al., 2023), and TimesNet (Wu et al., 2023). We evaluate all models on eight widely used datasets: Electricity, Weather, Illness, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). For detailed dataset descriptions, see Appendix  8.1.
@@
Methods ParallelTime S-Mamba iTransformer PatchTST DLinear TimesNet FEDFormer MOMENT GPT4TS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.189 0.158 0.210 0.168 0.219 0.149 0.198 0.176 0.237 0.172 0.220 0.217 0.296 0.154 0.209 0.162 0.212
192 0.189 0.232 0.203 0.252 0.211 0.255 0.194 0.241 0.220 0.282 0.219 0.261 0.276 0.336 0.197 0.248 0.204 0.248
336 0.242 0.273 0.258 0.292 0.260 0.292 0.245 0.282 0.265 0.319 0.280 0.306 0.339 0.380 0.246 0.285 0.254 0.286
720 0.323 0.331 0.328 0.340 0.332 0.341 0.314 0.334 0.333 0.362 0.365 0.359 0.403 0.428 0.315 0.336 0.326 0.337
ETTh1 96 0.365 0.398 0.395 0.422 0.407 0.428 0.370 0.399 0.375 0.399 0.384 0.402 0.376 0.419 0.387 0.410 0.376 0.397
192 0.399 0.415 0.427 0.443 0.427 0.443 0.413 0.421 0.405 0.416 0.436 0.429 0.420 0.448 0.410 0.426 0.416 0.418
336 0.385 0.414 0.462 0.469 0.456 0.463 0.422 0.436 0.439 0.443 0.491 0.469 0.459 0.465 0.422 0.437 0.442 0.433
720 0.420 0.443 0.522 0.518 0.468 0.472 0.447 0.466 0.472 0.490 0.521 0.500 0.506 0.507 0.454 0.472 0.477 0.456
ETTh2 96 0.262 0.328 0.298 0.356 0.298 0.357 0.274 0.336 0.289 0.353 0.340 0.374 0.358 0.397 0.288 0.345 0.285 0.342
192 0.322 0.368 0.372 0.399 0.377 0.406 0.339 0.379 0.383 0.418 0.402 0.414 0.429 0.439 0.349 0.386 0.354 0.389
336 0.312 0.370 0.402 0.432 0.424 0.440 0.329 0.380 0.448 0.465 0.452 0.452 0.496 0.487 0.369 0.408 0.373 0.407
720 0.399 0.434 0.419 0.449 0.438 0.462 0.379 0.422 0.605 0.551 0.462 0.468 0.463 0.474 0.403 0.439 0.406 0.441
ETTm1 96 0.284 0.337 0.309 0.361 0.313 0.367 0.290 0.342 0.299 0.343 0.338 0.375 0.379 0.419 0.293 0.349 0.292 0.346
192 0.329 0.366 0.345 0.384 0.348 0.385 0.332 0.369 0.335 0.365 0.374 0.387 0.426 0.441 0.326 0.368 0.332 0.372
336 0.365 0.391 0.375 0.403 0.377 0.403 0.366 0.392 0.369 0.386 0.410 0.411 0.445 0.459 0.352 0.384 0.366 0.394
720 0.424 0.430 0.435 0.440 0.438 0.438 0.416 0.420 0.425 0.421 0.478 0.450 0.543 0.490 0.405 0.416 0.417 0.421
ETTm2 96 0.162 0.252 0.177 0.270 0.179 0.273 0.165 0.255 0.167 0.269 0.187 0.267 0.203 0.287 0.170 0.260 0.173 0.262
192 0.218 0.291 0.229 0.305 0.242 0.315 0.220 0.292 0.224 0.303 0.249 0.309 0.269 0.328 0.227 0.297 0.229 0.301
336 0.276 0.327 0.281 0.338 0.291 0.345 0.274 0.329 0.281 0.342 0.321 0.351 0.325 0.366 0.275 0.328 0.286 0.341
720 0.356 0.380 0.371 0.392 0.377 0.398 0.362 0.385 0.397 0.421 0.408 0.403 0.421 0.415 0.363 0.387 0.378 0.401
Illness 24 1.166 0.657 1.918 0.847 1.960 0.952 1.319 0.754 2.215 1.081 2.317 0.934 3.228 1.260 2.728 1.114 2.063 0.881
36 1.293 0.727 2.006 0.944 2.264 0.978 1.430 0.834 1.963 0.963 1.972 0.920 2.679 1.080 2.669 1.092 1.868 0.892
48 1.399 0.772 2.080 0.898 2.266 1.042 1.553 0.815 2.130 1.024 2.238 0.940 2.622 1.078 2.728 1.098 1.790 0.884
60 1.615 0.844 2.414 1.094 2.541 1.108 1.470 0.788 2.368 1.096 2.027 0.928 2.857 1.157 2.883 1.126 1.979 0.957
Electricity 96 0.128 0.222 0.133 0.230 0.131 0.227 0.129 0.222 0.140 0.237 0.168 0.272 0.193 0.308 0.136 0.233 0.139 0.238
192 0.148 0.240 0.155 0.250 0.153 0.249 0.157 0.240 0.153 0.249 0.184 0.289 0.201 0.315 0.152 0.247 0.153 0.251
336 0.163 0.258 0.169 0.268 0.168 0.264 0.163 0.259 0.169 0.267 0.198 0.300 0.214 0.329 0.167 0.264 0.169 0.266
720 0.197 0.288 0.197 0.293 0.198 0.291 0.197 0.290 0.203 0.301 0.220 0.320 0.246 0.355 0.205 0.295 0.206 0.297
Traffic 96 0.349 0.231 0.354 0.252 0.350 0.257 0.360 0.249 0.410 0.282 0.593 0.321 0.587 0.366 0.391 0.282 0.388 0.282
192 0.371 0.240 0.373 0.260 0.387 0.276 0.379 0.256 0.423 0.287 0.617 0.336 0.604 0.373 0.404 0.287 0.407 0.290
336 0.388 0.250 0.390 0.265 0.407 0.289 0.392 0.264 0.436 0.296 0.629 0.336 0.621 0.383 0.414 0.292 0.412 0.294
720 0.429 0.274 0.430 0.288 0.433 0.297 0.432 0.286 0.466 0.315 0.640 0.350 0.626 0.382 0.450 0.310 0.450 0.312
@@
Table 1: The complete results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. Red: the best, Underline: the 2nd best.
@@
Table 2: Comparison of ParallelTime and PatchTST on the Traffic dataset. The table reports MSE, MAE, forward and backward (Fwd+Bwd) FLOPs (i.e., training FLOPs), and the number of parameters (#Params). Bold values indicate superior performance. ↓ indicates that lower values are better. The improvement percentages for ParallelTime over PatchTST are shown in parentheses.
MSE MAE Fwd+Bwd FLOPs #Params
Pred Len ParallelTime PatchTST ParallelTime PatchTST ParallelTime PatchTST ParallelTime PatchTST
96 0.349 (↓3.1%) 0.360 0.231 (↓7.2%) 0.249 25.2G (↓36%) 39.5G 614k (↓48%) 1194k
192 0.371 (↓2.1%) 0.379 0.240 (↓6.3%) 0.256 25.2G (↓37%) 40.5G 651k (↓67%) 1980k
336 0.388 (↓1.0%) 0.392 0.250 (↓5.3%) 0.264 25.3G (↓39%) 42.1G 707k (↓77%) 3160k
720 0.429 (↓0.7%) 0.432 0.274 (↓4.2%) 0.286 25.5G (↓44%) 46.3G 855k (↓86%) 6306k
@@ 4.2 Main Results
Comprehensive forecasting results are listed in Table 1, with the best performance highlighted in red and the second best underlined. All model results are from (Goswami et al., 2024), except S-mamba and iTransformer, which we trained due to unavailable results for window size 512. A lower Mean Squared Error (MSE) and Mean Absolute Error (MAE) indicate more accurate predictions. Our proposed model, ParallelTime, demonstrates exceptional performance across a diverse set of datasets and prediction horizons, consistently outperforming a range of state-of-the-art models, ParallelTime achieves the best forecasting accuracy in a significant number of scenarios, particularly excelling in datasets such as Weather, ETTh1, ETTh2, ETTm2, Electricity, Traffic, and Illness.